OpenStudy (anonymous):

Solve (x-9)^2=1 Can someone show me the steps to solve this?

OpenStudy (experimentx):

(x-9)^2 = 1^2 which means (x-9) = +1 or (x-9) = -1

OpenStudy (diyadiya):

\[ (x-9)^2=1\] taking square root on both sides \[(x-9)= \pm 1\] \[(x-9)=1 \tag{1}\]\[(x-9)=-1 \tag{2}\] Now solve for x